Earth absorb 10^22 joule of energy every day from sun. What will % Inc in the mass of earth due to this

Grade:12th pass

1 Answers

Khimraj
5 years ago
The sun strikes the Earth with 1022 J of energy every day. Using  E = mc2 
m = 1022/9*1016 = 1.11*105 kg